# Identifying and Addressing Delusions for Target-Directed Decision Making

### v3

@@ 3.2 Estimator Delusions
This type of delusion appears when an estimator mis-evaluates the favorability of a G.3.1 target. If effective learning rules are present, this can still be caused by the lack of training data sampling procedure (e.g., existence of relabeled transitions with G.3.1 targets in HER). In Fig. 2 a), an example is visualized. Note that the resulting E.3.2.1 behaviors, i.e. those chasing selected G.3.1 targets, can be potentially catastrophic if the G.3.1 targets are beyond safety constraints.
@@ 4 Addressing Delusions: Strategies Applied to Hindsight Relabeling
Identifying delusions enables mitigation strategies. We propose general solutions for existing target-directed agents, coming from a spectrum of training procedures. Given the popularity of HER-based training and the generality behind hindsight relabeling, we provide examples of applying our strategies to address delusions in HER-based training, demonstrating empirical benefits.
Assume we are dealing with a hypothetical target-directed agent with a dual-component architecture as in Fig. 1, learning both components exclusively from hindsight-relabeled transitions. The generator proposes targets corresponding to potentially distant states, and the estimator learns the favorability of targets (via an oracle or more complex methods) by punishing the unachieved or unsafe targets present in the training data. To address delusions, we combine update rules and training data distribution. By assuming this framework, we skip discussing update rules, as they depend on the agent’s design.
@@ 4.1 Assistive Strategies for Addressing Delusions
We first introduce two strategies to address delusions, which are then materialized as two hindsight relabeling strategies for HER. Later, we compare the ideas behind the new strategies with existing ones, and study how the two groups can be combined to create mixtures for estimators and hybrids for the generator-estimator duo.
The following strategies, which creates the training data from experienced episodes, seek to expand the support of the training data distribution, to include those source-target combinations involving G.3.1 and G.3.1.1 targets, respectively. With the assumed update rules, this extra exposure is expected to address E.3.2.1 and E.3.2.2 estimator delusions.
@@
In Zhao et al. (2024), the authors identified a delusional behavior resulting from, E.3.2.1 delusions in the language of this paper, and they proposed to train the estimator additionally with candidate targets proposed by the generator. With HER, we can transform this auxiliary loss into a Just-In-Time (JIT) HER strategy that, when a transition is sampled for training, relabels it with a conditionally generated target by the generator. “generate” is particularly effective for training estimators, as the estimators will get exposed to all kinds of problematic targets (that the generator could offer [^1] When a generator learns to generate mostly viable targets, “generate” will be primarily effective at dealing with the E.3.2 and E.3.2.1 (G.3.1s are unlikely to be eradicated). “generate” requires the use of the generator, thus it incurs additional computational burden.
Adapting “generate” beyond HER should be straightforward. Whenever a target is involved (during estimator learning), we can replace it with generated targets.
4.1.2 pertask”: Let Estimators Learn About Experienced Targets
The second strategy, namely “pertask”, is to expose the estimator to targets that were experienced before, s.t. it could figure out that some targets are not favorable from the current state.
We materialize “pertask” into an assistive atomic relabeling strategy that relabels transitions with observations from the same training task, sampled across the entire memory. “pertask” works on a task level, creating more diverse source-target pairs than trajectory-level strategies like “future” & “episode”, which bias the training data distribution toward shorter-distance source-target pairs. This approach additionally brings exposure to the estimator to learn against E.3.2.2 delusions and to the long-distance source-target pairs, thus mitigating E.3.2 caused by short trajectories.
@@ 4.2 Mixtures
Mixtures of multiple relabeling strategies can also be extended to other training procedures, and can also be implemented by independent training procedures based on different sources.
@@ 4.3 Hybrid Strategies - A $2$-Slotted Approach
While an estimator needs the exposure to problematic targets to counteract delusions, a generator benefits from learning to generate useful targets, and should avoid exposure to problematic targets. Thus, generators and estimators often have conflicting needs during training.
@@ 5 Related Works
Hindsight Relabeling (Andrychowicz et al., 2017) is crucial for enhancing sample efficiency in target-directed RL, as it enables learning from failed experiences (Dai et al., 2021). From the perspective of source-target pair distribution, from which we similarly developed this paper, improvements were proposed, including a divergence maximization approach (Zhang & Stadie, 2022), addressing distributional change (Bai et al., 2023), and multistep relabeling strategies (Yang et al., 2021b), etc. Kuang et al. (2020) proposed to prioritize rarely-seen achieved transitions by utilizing the density distribution. HER’s success was centered on its performance gains via sample efficiency (in non-delusional cases), around which most follow-up works revolved as well.
While, performance deficiencies of target-directed frameworks can be resulted from multiple factors, which include delusions. Shams & Fevens (2022) studied the performance of atomic strategies from the view of sample efficiency, without looking into the failure modes. Deshpande et al. (2018) detailed experimental techniques in sparse reward settings using “future”. In (Yang et al., 2021a), a mixture strategy similar to “generate” improved non-delusional estimations, though the impact on delusions was not explored, possibly because of the lack of an appropriate RL environment. Nasiriany et al. (2019) used a mixture of up to $3$ atomic strategies for producing effective training data in a single-task setting, while its contribution to addressing delusions was not investigated. The performance of existing target-directed agents are often limited by their exclusive reliance on “future” or “episode(He et al., 2020), whose delusions this paper intends to address.
Delusions in RL. Delusions can stem from various other improper designs. For example, Lu et al. (2018) identified delusions caused by the limitations of function approximators for greedy policies in modelfree RL; Often accompanying hallucinations, delusions can arise in non-target-directed frameworks as well. In an offline planning framework such as Dyna, state hallucinations can irreversibly destabilize value estimations (Jafferjee et al., 2020). In Di Langosco et al. (2022); Shah et al. (2022), the authors classified goal misgeneralization, a class of delusional behaviors with which an agent competently pursues an undesired target leading in novel test situations. Zhao et al. (2024) gave first examples of delusional behaviors caused by improper hindsight relabeling.
@@ 6.3 Methods
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitest sequence is chosen to be the target, and used to condition a lower-level policy. LEAP is prone to delusions, since one problematic subgoal and destroy a whole sequence (Nasiriany et al., 2019).
@@ 6.5 Skipper on SSM (Set 1/4)
6.6 Summary of Experiments & Empirical Guidelines
@@ 7 Empirical Guidelines
The following steps can be applied to target-directed methods beyond HER:
@@ 8 Conclusion & Future Work
We identified the types and the causes of delusional behaviors, a shared failure mode of target-directed agents. When applying our strategies on HER, the resulting hybrid hindsight relabeling strategies flexibly satisfy the needs of the generator and the estimator, while avoiding the naïveté in training processes that could lead to more delusional behaviors and catastrophic safety breaches.
@@ Appendix A Summaries
Inappropriate Update Rules Some update rules do not lead to punishment of the favorability of unreachable targets Adjust the learning rules
Bad Training Data Distributions the collected experience was not organized in a way that produced sufficiently diverse source-target distributions to make the agent realize, via appropriate update rules, that wrong beliefs are wrong Adjust how training data distribution is crafted (minding G.3.1 and G.3.1.1).

### v4

@@ 3.2 Estimator’s Role in Delusional Behaviors: Bad Evaluation
This type of delusion appears when an estimator misevaluates the favorability of a G.3.1 target. If effective learning rules are present, this can still be caused by the lack of necessary training data (e.g., relabeled transitions with G.3.1 targets in buffer). In Fig. 2 a), an example is visualized. Note that the resulting E.3.2.1 behaviors, i.e. those chasing selected G.3.1 targets, can be potentially catastrophic if the G.3.1 targets are beyond safety constraints.
@@ 4 Addressing Delusions: Strategies Applied to Hindsight Relabeling
Identifying delusions enables mitigation strategies. We develop methodologies applicable to general target-directed agents coming from a spectrum of training procedures. Then, we provide examples on how we can materialize such methodologies in agents trained with HER, because of its popularity and the generality behind hindsight relabeling (and source-target pairs).
Assume we are dealing with a hypothetical target-directed agent with a dual-component architecture as in Fig. 1, learning both components exclusively from hindsight-relabeled transitions. The generator proposes targets corresponding to potentially distant states, and the estimator learns the favorability of targets by (continuously) punishing the unachieved or unsafe targets present in the training data. As discussed earlier, to address delusions, we need effective update rules and training data distribution. By assuming this framework, we skip discussing update rules, as they depend on the agents’ specific designs. Our emphasis on the training data reflects the fact that most existing RL agents only learn from experienced data, while addressing delusions requires learning from targets that can never be experienced. This corresponds to a discrepancy between target-directed agents’ behaviors and training: at decision-time, targets outside experience can be proposed; while during training, only experienced targets are learned from.
@@ 4.1 Assistive Strategies for Addressing Delusions
We first introduce two strategies to improve training data distributions, which are then materialized as two hindsight relabeling strategies for HER. Later, we compare the ideas behind the proposed strategies with existing ones, and study how the two groups can be combined to create mixtures for estimators and hybrids for the generator-estimator duo.
The following strategies seek to expand the support of the training data distribution, to include those source-target combinations that the agent could never experience, i.e. those involving G.3.1 and G.3.1.1 targets. With the assumed update rules, this extra exposure is expected to address E.3.2.1 and E.3.2.2 estimator delusions.
@@
In Zhao et al. (2024), the authors identified a delusional behavior resulting from E.3.2.1 delusions, in the language of this paper, and they proposed to train the estimator additionally with candidate targets proposed by the generator. With HER, we can transform this auxiliary loss into a Just-In-Time (JIT) HER strategy that, when a transition is sampled for training, relabels it with a conditionally generated target by the generator. We can expect “generate” to be effective for training estimators, as the estimators will get exposed to all kinds of problematic targets that the generator could offer: When a generator learns to generate mostly viable targets, “generate” will be primarily effective at dealing with the E.3.2 and E.3.2.1 (G.3.1s are unlikely to be eradicated). “generate” requires the use of the generator, thus it incurs additional computational burden, depending on the complexity of target generation processes.
Adapting “generate” beyond HER should be straightforward: whenever a target is involved during estimator learning, we may replace it with generated targets.
4.1.2 pertask”: Let Estimators Learn About Experienced Targets Outside the Episode
The second strategy, namely “pertask”, is to expose the estimator to targets that were experienced before, s.t. it could figure out that some previously achieved targets are not achievable from the current state.
We materialize “pertask” into an assistive atomic relabeling strategy that relabels transitions with observations from the same training task, sampled across the entire memory. “pertask” works on a task level, creating more diverse source-target pairs than trajectory-level strategies like “future” & “episode”, which bias the training data distribution toward shorter-distance source-target pairs. This approach additionally brings exposure to the estimator to learn against E.3.2.2 delusions and to the long-distance source-target pairs against E.3.2 caused by short trajectories.
@@ 4.2 Mixtures
Mixtures of multiple relabeling strategies can also be extended to other training procedures, and can also be implemented by independent training losses based on different sources.
@@ 4.3 Hybrid Strategies - A $2$-Slotted Approach
While an estimator needs the exposure to problematic targets to counteract delusions, a generator benefits from learning to generate only useful targets, and should avoid exposure to problematic targets. Thus, generators and estimators often have conflicting needs during training.
@@ 5 Related Works
Hindsight Relabeling is crucial for enhancing sample efficiency in goal-directed RL (Andrychowicz et al., 2017), as it enables learning from failed experiences (Dai et al., 2021). From the perspective of source-target pairs, improvements were proposed, including divergence maximization (Zhang & Stadie, 2022), addressing distributional change (Bai et al., 2023), multistep relabeling strategies (Yang et al., 2021b), prioritizing rarely-seen achieved transitions (Kuang et al., 2020), etc.
HER’s success was centered on its performance gains via sample efficiency (in non-delusional cases), around which most follow-up works revolved as well. In reality, performance deficiencies of target-directed frameworks can be resulted from multiple factors, one of which is delusions. Shams & Fevens (2022) studied the performance of atomic strategies from the view of sample efficiency, without looking into the failure modes. Deshpande et al. (2018) detailed experimental techniques in sparse reward settings using “future”. In (Yang et al., 2021a), a mixture strategy similar to “generate” improved non-delusional estimations, though the impact on delusions was not explored, possibly because of the lack of an appropriate environment. Nasiriany et al. (2019) used a mixture of up to $3$ atomic strategies for producing effective training data in a single-task setting, while the contributions of the mixture to addressing delusions were not investigated. It is widely known that performance of existing HER-trained agents are often limited by their exclusive reliance on “future” or “episode(He et al., 2020), whose delusions this paper intends to address.
Delusions in RL can stem from various other improper designs. For example, Lu et al. (2018) identified delusions caused by the limitations of function approximators for greedy policies in modelfree RL; Often accompanying hallucinations, delusions can arise in non-target-directed frameworks as well. In an offline planning framework such as Dyna, state hallucinations can irreversibly destabilize value estimations (Jafferjee et al., 2020). In Di Langosco et al. (2022), the authors classified goal misgeneralization, a class of delusional behaviors with which an agent competently pursues an undesired target leading in novel test situations. Zhao et al. (2024) gave first examples of delusional behaviors caused by improper hindsight relabeling. The perspective of delusions can often be overlooked because of the lack of estimators in target-directed methods, as discussions regarding hallucinations get magnified.
@@ 6.3 Methods
LEAP (in Appendix): Instead of VI, LEAP uses an evolutionary algorithm to evolve sequences of subgoals leading to the task goal, with distances as fitness. The immediate subgoal of the elitest sequence is chosen to be the target, and used to condition a lower-level policy. LEAP is prone to delusions, since one problematic subgoal can destroy a whole sequence (Nasiriany et al., 2019).
@@ 6.5 Skipper on SSM (Set 1/4)
6.6 Summary of Experiments
@@ 7 Empirical Guidelines
The following steps can be applied to target-directed methods (applicable also beyond HER):
@@ 8 Conclusion & Future Work
We investigated the causes of delusional behaviors, a shared failure mode of target-directed agents. When applying our mitigation strategies on HER, the resulting hybrid hindsight relabeling strategies flexibly satisfy the needs of the generator and the estimator, granting the agents the ability to address delusions autonomously and preemptively avoid delusional behaviors.
@@ Appendix A Summaries
Inappropriate Update Rules Some update rules do not lead to punishment of the favorability of unachieved targets Adjust the learning rules to continuously punish the unachieved targets
Bad Training Data Distributions the collected experience was not organized in a way that produced sufficiently diverse source-target distributions to make the agent realize, via appropriate update rules, that wrong beliefs are wrong Augment the training data with source-target pairs involving targets that could participate in planning, but will never be experienced, i.e. G.3.1 and G.3.1.1